OpenStudy (anonymous):

d^2y/dx^2+4dy/dx+13y=2cos2x find the complementary function and particular integral.Hence write down the full general solutions.

zepdrix (zepdrix):

eyyyy :) What part are you having trouble with?

zepdrix (zepdrix):

If we look for just the solutions to the homogeneous first,\[\large\rm y''+4y'+13y=0\]We get a characteristic equation that looks like this,\[\large\rm r^2+4r+13=0\]Understand that part? :)

OpenStudy (anonymous):

ys i know but how to solve when right having 2cos2x

zepdrix (zepdrix):

Were you able to find the homogeneous solution? It should give you something like this:\[\large\rm y_h=c_1e^{-2x}\cos(3x)+c_2e^{-2x}\sin(3x)\]Then to find the particular solution, we assume that it has the form of the right side, more generally though, sines and cosines. \[\large\rm y_p=A\sin(2x)+B\cos(2x)\]

zepdrix (zepdrix):

And then we take a couple derivatives to find our \(\large\rm y'_p\) and \(\large\rm y''_p\) and then plug all of that back into the orginal differential equation.

zepdrix (zepdrix):

What do you think? :d Too confusing?

OpenStudy (anonymous):

oh i knew now! thx

zepdrix (zepdrix):

cool! :)